Find the general solution of the following differential equations
12y"-5y' - 2y = 0
1
Expert's answer
2021-09-22T16:54:46-0400

Corresponding (auxiliary) equation


"12r^2-5r-2=0"

"d=(-5)^2-4(12)(-2)=121"

"r=\dfrac{5\pm\sqrt{121}}{2(12)}=\dfrac{5\pm11}{24}"

"r_1=\dfrac{5-11}{24}=-\dfrac{1}{4},r_2=\dfrac{5+11}{24}=\dfrac{2}{3}"

The general solution of the given differential equation is


"y(t)=c_1e^{-{1 \over 4}t}+c_2e^{-{2 \over 3}t}"